# Transfer Learning for High-dimensional Quantile Regression with Distribution Shift

### v1

@@ 2.2 Theory on Estimation Error
Now we study the non-asymptotic convergence rate of $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ obtained by the transfer approach (4) with transferable set $\mathcal{C}$, where $\mathcal{C}$ is an arbitrary prespecified set for theoretical generality. Under the high-dimensional sparse setting, we characterize the geometric structure of $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ and $\{\widehat{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}\}_{k=1}^{K}$ through the restrictive sets (Belloni and Chernozhukov,, 2011; Negahban et al.,, 2012; Wang and He,, 2024), which are denoted by
@@
for $\widehat{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}$ with $\ell_{1}$-sparsity: $\Gamma_{W}^{(k)}=\{v\in\mathbb{R}^{p}:\|v_{S_{a}^{(k)c}}\|_{1}\leq\|v_{S_{a}^{(k)}}\|_{1}+2\|\boldsymbol{\delta}^{(k)}_{S_{a}^{(k)c}}\|_{1}\},$
where $S=\{j:\boldsymbol{\beta}^{*}_{j}\neq 0,1\leq j\leq p\},S_{a}^{(k)}=\{j:|\boldsymbol{\delta}^{(k)}_{j}|>a,1\leq j\leq p\}$ with a threshold $a>0$. The sets $\{\Gamma_{W}^{(k)}\}_{k=1}^{K}$ depends on $a$, but we omit the dependence in notation for simplicity. $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ and $\{\widehat{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}\}_{k=1}^{K}$ will be shown to lie in the sets $\Gamma_{H}$ and $\{\Gamma_{W}^{(k)}\}_{k=1}^{K}$ with high probability respectively in the supplementary materials. Conditions below constitute a set of basic assumptions for establishing the statistical properties, where the index $k$ refers to each study $k\in\{0,\dots,K\}$.
@@
The conditional density value at the $\tau$-th quantile $f_{i}^{(k)}=f_{i}^{(k)}(0|\boldsymbol{x}_{i}^{(k)})$ is positive, continuously differentiable, and uniformly upper bounded. Also, there exist a universal positive constant $b_{0}$ such that the derivatives $f_{i}^{(k)}\text{}^{\prime}(t|\boldsymbol{x}_{i}^{(k)})$ exist and are uniformly bounded for $|t|\leq b_{0}$.
Condition 2.
@@
Each covariate $\boldsymbol{x}_{i}^{(k)}$ is a mean-zero sub-Gaussian random vector with sub-Gaussian norm uniformly bounded by a positive constant $m_{g}$.
Condition 4.
There exists $a_{k}>0$ and corresponding index set $S_{a_{k}}^{(k)}$ such that $\boldsymbol{\delta}^{(k)}=\boldsymbol{\delta}^{(k)}_{S_{a_{k}}^{(k)}}+\boldsymbol{\delta}^{(k)}_{S_{a_{k}}^{(k)c}}$ with $\|\boldsymbol{\delta}^{(k)}_{S_{a_{k}}^{(k)}}\|_{0}\leq s_{k},\|\boldsymbol{\delta}^{(k)}_{S_{a_{k}}^{(k)c}}\|_{1}\leq\eta_{k}$.
Local boundness and smoothness of conditional density functions in Condition 1 are very familiar in QR literatures. Here a uniform lower bound of $f_{i}^{(k)}$ is not assumed, since we want to reveal the effect of residual shift in the error bounds. Condition 2 assumes a modified version of restricted eignevalue condition (RSC) with incorporating the $\tau$-th density $\mathbb{E}[f_{i}^{(k)}]$. The reason for placing $\mathbb{E}[f_{i}^{(k)}]$ on the right-hand side is to technically derive error bounds involving each study’s $\tau$-th density. Note that if the study model is homogeneous, then this condition will return to the original RSC form. Condition 3 needs covariates following sub-Gaussian distributions, which relaxes the former requirement of bounded design (Tian and Feng,, 2023; Huang et al.,, 2023). As for Condition 4, we impose an approximate sparse structure on $\boldsymbol{\delta}^{(k)}$ with the sparse support $S_{a_{k}}^{(k)}$ to be different across source studies, which is also similarly assumed in Zhang and Zhu, (2023) and Huang et al., (2023).
@@
Assume that Conditions 1-4 is satisfied. Define $n_{\mathcal{C}}=n_{0}+\sum_{k\in\mathcal{C}}n_{k}$. Let $\widehat{\boldsymbol{\beta}}$ is obtained from (4) with transferable set $\mathcal{C}$. For each study $k\in\{0\}\cup\mathcal{C}$, let
$\lambda_{\boldsymbol{\beta}}=c_{\lambda}\sqrt{n_{\mathcal{C}}\log p},\qquad\lambda_{k}=c_{\lambda}n_{k}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}},$
where $c_{\lambda}$ is a large enough constant. Let $r_{k}=s_{k}\sqrt{\log p/(n_{k}\wedge n_{0})}+\eta_{k}$, if
@@
The assumption (6) is a technical requirement to bound the infinite norms of subgradients in our proofs. Note that $r_{k}$ is actually the $\ell_{1}$-error bound of each $\widehat{\boldsymbol{\delta}}^{(k)}$ under soft sparsity. A sufficient condition for (6) is to let $s^{6}(\log p)^{2}=o(n_{k})$ and $r_{k}=o(n_{k}^{-1/4})$ for each $k\in\mathcal{C}$, which is not harsh in the high-dimensional setting. Compared to Theorem 1 in Li et al., (2024) under the similar transfer framework, we relaxed their target size assumption $|\mathcal{C}|n_{0}\ll n_{\mathcal{C}}$, and our choice for $\lambda_{\boldsymbol{\beta}}$ frees the dependence on the unknown parameters $s$ and $h_{1}$. Moreover, our analysis allows the size of transferable set $|\mathcal{C}|$ goes to infinity.
@@
The $\ell_{2}/\ell_{1}$-error bounds of $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ in (7)/(8) contain two parts. The first part can be seen as the inherent estimation error $\sqrt{s\log p/n_{\mathcal{C}}}$ accompanied by the variance term $\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[f_{i}^{(k)}]$ from the residual shift. The second part represents the bias term involving each contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}$ from source’s parameter shift weighted by the sample size $n_{k}$, and also includes the variance term. The error bounds reveal that to improve the transfer performance, source studies need to possess small $\|\boldsymbol{\delta}^{(k)}\|_{1}$ as well as large $\mathbb{E}[f_{i}^{(k)}]$, strongly confirming the rationality of our screening criteria (3). Due to the weighted summation form, the optimal informative level of $\|\boldsymbol{\delta}^{(k)}\|_{1}$ and $\mathbb{E}[f_{i}^{(k)}]$ is nearly impossible to derive, since the impact of each source depends on its sample size proportion $n_{k}/n_{\mathcal{C}}$ and other sources’ shift level, let alone the implicit dynamic choice of set $\mathcal{C}$. That’s why we offer the compromised universal levels $h_{1},h_{2}$ in our transferable set $\mathcal{C}_{h}$. For better understanding of our Theorem 1, here we give several quantitative explanations under some simplified settings, taking the $\ell_{2}$ error bound (7) as the example.
  • (Parameter Shift) If all $\tau$-th densities $\mathbb{E}[f_{i}^{(k)}]$ are neglected, then the result (7) can be simplified to
    $\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}\|_{2}\lesssim\sqrt{\frac{s\log p}{n_{\mathcal{C}}}}+\sqrt{\sum_{k\in\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}\|\boldsymbol{\delta}^{(k)}\|_{1}}.$
@@
    Parallel to the previous transfer theory, the first term $\sqrt{s\log p/n_{\mathcal{C}}}$ enjoys a sharper rate than the single-task $\ell_{2}$-error $\sqrt{s\log p/n_{0}}$ (Belloni and Chernozhukov,, 2011), which indeed reflects the superiority of transfer learning. The second term is a weighted bias term by parameter shift, and if $n_{0}=O(n_{k}),k\in\mathcal{C}$, this term can be simplified to $(\log p/n_{0})^{1/4}\sqrt{\max_{k\in\mathcal{C}}\|\boldsymbol{\delta}^{(k)}\|_{1}}$, corresponding to the result in Theorem 1 of Li et al., (2024). When the contrast magnitude $\max_{k\in\mathcal{C}}\|\boldsymbol{\delta}^{(k)}\|_{1}=o(s\sqrt{\log p/n_{0}})$, i.e., smaller than the order of single-task $\ell_{1}$-error, the transfer estimator will improve the target task performance.
@@
  • (Residual Shift) To our knowledge, this non-asymptotic result containing $\tau$-th density $\mathbb{E}[f_{i}^{(k)}]$ is new to the transfer literature, since we keep the density items along our proving process. For the first term, we can give a simple example to show that why residual shift could impact the transfer performance. Comparing the single-task $\ell_{2}$-error to the error (7) with only one source study, if the source study has huge residual shift such that $\mathbb{E}[f_{i}^{(1)}]<\frac{\sqrt{n_{0}}(\sqrt{n_{0}+n_{1}}-\sqrt{n_{0}})}{n_{1}}\mathbb{E}[f_{i}^{(0)}]$, then incorporating this source study will lead to negative transfer. In the second term, for simplicity assuming $n_{0}=O(n_{k}),k\in\mathcal{C}$, the error changes to
    $\sqrt{\sqrt{\frac{\log p}{n_{0}}}\cdot\frac{\sum_{k\in\mathcal{C}}n_{k}\|\boldsymbol{\delta}^{(k)}\|_{1}}{\sum_{k\in\{0\}\cup\mathcal{C}}n_{k}\mathbb{E}[f_{i}^{(k)}]}}.$
    The first part $\sqrt{\log p/n_{0}}$ is inherited from the target study, since we need the target information to estimate each $\boldsymbol{\delta}^{(k)}$. The second part clearly reveals the effect of both parameter/residual shift weighted by sample sizes, which again corroborates the validity of the construction of our transferable set.
@@
  • (Covariate Shift) Our framework can relax the elusive conditions about the covariance matrices across studies (Assumption 4 in Tian and Feng, (2023), Condition 4 in Zhang and Zhu, (2023), Assumption 4 in Huang et al., (2023)). In contrast, the two-step transfer framework may amplify the error bounds by a factor $C_{\boldsymbol{\Sigma}}$, which may can diverge when the covariance matrices $\boldsymbol{\Sigma}^{(k)},k\in\mathcal{C}$ are dissimilar to $\boldsymbol{\Sigma}^{(0)}$. To put it succinctly, our results neither assume the similarity of $\boldsymbol{\Sigma}^{(k)}$, nor involve the factor $C_{\boldsymbol{\Sigma}}$ mentioned above, hence achieving better performance under covariate shift.
@@
Under Conditions 1-4 and assumptions in Theorem 1, take the oracle transferable set $\mathcal{C}=\mathcal{C}_{h}$ and suppose $n_{0}=O(n_{k}),k\in\mathcal{C}_{h}$. Then for any constant $\varepsilon>0$, with probability at least $1-\varepsilon-(|\mathcal{C}_{h}|+1)A_{1}\exp(-A_{2}\log p)$, we can obtain that
$\|\widehat{\boldsymbol{\beta}}_{ora}-\boldsymbol{\beta^{*}}\|_{2}\lesssim\sqrt{\frac{s\log p}{n_{\mathcal{C}_{h}}}}h_{2}+(\frac{\log p}{n_{0}})^{\frac{1}{4}}\sqrt{h_{1}h_{2}},$
As shown in Corollary 1, the terms $h_{1}$ and $h_{2}$ regulate the degree of parameter and residual shift, respectively, striking a balance between performance improvement and the risk of negative transfer. For the first term, a sufficient condition to avoid negative transfer is to ensure that $h_{2}=o(\sqrt{n_{\mathcal{C}_{h}}/n_{0}})$, which indicates that the minimum $\tau$-th density in the oracle transferable set should exceed $\sqrt{n_{0}/n_{\mathcal{C}_{h}}}\mathbb{E}[f_{i}^{(0)}]$. The second term highlights the nuanced relationship between the parameter shift $h_{1}$ (measured by $\max_{k\in\mathcal{C}_{h}}\|\boldsymbol{\delta}^{(k)}\|_{1}$) and the residual shift $h_{2}$ (measured by $\max_{k\in\mathcal{C}_{h}}\mathbb{E}[f_{i}^{(0)}]/\mathbb{E}[f_{i}^{(k)}]$), with a sufficient condition being $h_{1}h_{2}=o(s\sqrt{\log p/n_{0}})$. This generalizes the previous result $(\log p/n_{0})^{1/4}\sqrt{h_{1}}$ in Li et al., (2024), allowing for better error control by managing both parameter and residual shift.

### v2

@@ 2.2 Theory on Estimation Error
Now we study the non-asymptotic convergence rate of $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ obtained by the transfer approach (3) with transferable set $\mathcal{C}$, where $\mathcal{C}$ is an arbitrary prespecified set for theoretical generality. Under the high-dimensional sparse setting, we characterize the geometric structure of $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ and $\{\widehat{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}\}_{k=1}^{K}$ through restrictive sets [belloni2011l1, wang2024analysis], denoted by
@@
for $\widehat{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}$ with $\ell_{1}$-sparsity: $\Gamma_{W}^{(k)}=\{v\in\mathbb{R}^{p}:\|v_{S_{k}^{c}}\|_{1}\leq\|v_{S_{k}}\|_{1}+2\|\boldsymbol{\delta}^{(k)}_{S_{k}^{c}}\|_{1}\},$
where $S=\{j:\boldsymbol{\beta}^{*}_{j}\neq 0,1\leq j\leq p\}$, $S_{k}=\{j:|\boldsymbol{\delta}^{(k)}_{j}|>a_{k},1\leq j\leq p\}$ with a threshold $a_{k}>0$. Furthermore, for each contrast vector $\boldsymbol{\delta}^{(k)}$, we denote $s_{k}=\|\boldsymbol{\delta}^{(k)}_{S_{k}}\|_{0},\eta_{k}=\|\boldsymbol{\delta}^{(k)}_{S_{k}^{c}}\|_{1}$. As noted in [wang2024analysis], $a_{k}$ controls the approximate sparsity level of the contrast vector $\boldsymbol{\delta}^{(k)}$. Although the exact value of each $a_{k}$ is not specified, the constraints on $s_{k}$ and $\eta_{k}$ in our subsequent theorem impose implicit requirements on $a_{k}$. For simplicity, we omit the dependence on $a_{k}$ in the notations $\Gamma_{W}^{(k)},s_{k},\eta_{k}$. We introduce three conditions below, where the index $k$ refers to each study $k\in\{0\}\cup\mathcal{C}$.
@@
Condition 2.
Each covariate $\boldsymbol{x}_{i}^{(k)}$ is a mean-zero $m_{g}$-sub-Gaussian random vector, where the sub-Gaussian norm $m_{g}$ is a universal positive constant.
@@
The conditional density function $f^{(k)}(\cdot|\boldsymbol{x}_{i}^{(k)})$ is positive, continuously differentiable, and uniformly upper bounded. Also, there exist a universal positive constant $b_{0}$ such that the derivative $f_{i}^{(k)^{\prime}}(t|\boldsymbol{x}_{i}^{(k)})$ is uniformly bounded for $|t|\leq b_{0}$.
Condition 1 assumes a modified version of restricted eigenvalue condition (RSC) with incorporating the $\tau$-th density $\mathbb{E}[f_{i}^{(k)}]$. The reason for placing $\mathbb{E}[f_{i}^{(k)}]$ on the right-hand side is to technically derive error bounds involving each study’s $\tau$-th density. Note that if the study model is homoscedastic, then this condition will return to the original RSC form. Condition 2 needs covariates following sub-Gaussian distributions, which relaxes the former requirement of bounded design [tian2022transfer, huang2023estimation]. Condition 3 requires the boundness and smoothness of conditional density functions. Here a uniform lower bound of $f_{i}^{(k)}$ is not assumed, since we want to reveal the effect of residual shift in the error bounds. All three conditions are very common in the high-dimensional QR literature [belloni2011l1, wang2024analysis].
@@
Assume Conditions 1-3 are satisfied. Define $n_{\mathcal{C}}=n_{0}+\sum_{k\in\mathcal{C}}n_{k}$. Let $\widehat{\boldsymbol{\beta}}$ be obtained from the transfer framework (3) with the transferable set $\mathcal{C}$. For each study $k\in\{0\}\cup\mathcal{C}$, let $\lambda_{\boldsymbol{\beta}}=c_{\lambda}\sqrt{n_{\mathcal{C}}\log p},\lambda_{k}=c_{\lambda}n_{k}\sqrt{\log p/(n_{k}\wedge n_{0})}$, where $c_{\lambda}$ is a large enough constant. Let $r_{k}=s_{k}\sqrt{\log p/(n_{k}\wedge n_{0})}+\eta_{k}$, if
@@
The assumption (4) is a technical requirement for parameter growth, where $r_{k}$ is actually the $\ell_{1}$-error bound of each $\widehat{\boldsymbol{\delta}}^{(k)}$ under soft sparsity. A sufficient condition for (4) is to let $s^{6}(\log p)^{2}=o(n_{k})$ and $r_{k}=o(n_{k}^{-1/4})$ for each $k\in\mathcal{C}$, which is not harsh in the high-dimensional setting. Compared to Theorem 1 in [li2023estimation] under the similar transfer framework, we relaxed their target size assumption $|\mathcal{C}|n_{0}\ll n_{\mathcal{C}}$, and our choice for $\lambda_{\boldsymbol{\beta}}$ frees the dependence on the unknown parameters $s$ and $h_{1}$. Moreover, our analysis allows the size of transferable set $|\mathcal{C}|$ goes to infinity.
@@
The error bounds of $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ in (5)/(6) contain two parts. The first part can be seen as the inherent estimation error $\sqrt{s\log p/n_{\mathcal{C}}}$ weighted by the variance term $\sum_{k\in\{0\}\cup\mathcal{C}}n_{k}\mathbb{E}[f_{i}^{(k)}]$ from residual shift. The second part represents the bias term involving each contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}$ from parameter shift weighted by the variance term. The error bounds reveal that to improve the transfer performance, source studies need to possess small $\|\boldsymbol{\delta}^{(k)}\|_{1}$ as well as large $n_{k}\mathbb{E}[f_{i}^{(k)}]$, strongly confirming the rationality of our screening criteria (2). Here we give some quantitative explanations under simplified settings, taking the $\ell_{2}$-error bound (5) as the example.
@@
  • (Parameter Shift) If the $\tau$-th densities $\mathbb{E}[f_{i}^{(k)}]$ are neglected and $n_{0}=O(n_{k}),k\in\mathcal{C}$, then the error bound can be simplified to $\sqrt{s\log p/n_{\mathcal{C}}}+(\log p/n_{0})^{1/4}\sqrt{\max_{k\in\mathcal{C}}\|\boldsymbol{\delta}^{(k)}\|_{1}}$. Parallel to the previous transfer theory, the first term $\sqrt{s\log p/n_{\mathcal{C}}}$ enjoys a sharper rate than the single-task $\ell_{2}$-error $\sqrt{s\log p/n_{0}}$ [belloni2011l1, wang2024analysis], which indeed reflects the superiority of transfer learning. The second term is a bias term by parameter shift, corresponding to the result in Theorem 1 of [li2023estimation].
@@
  • (Residual Shift) To our knowledge, this non-asymptotic result containing $\tau$-th density $\mathbb{E}[f_{i}^{(k)}]$ is new to the transfer literature, since we keep the density items along our proving process. For both terms, we can observe clearly from the denominator that the contribution of each study is represented by the product $n_{k}\mathbb{E}[f_{i}^{(k)}]$, which reflects the information quality around the objective quantile and reinforces the reasonability of our transferability level $h_{2}$ defined in (2). For example, consider the case with only one source study that exhibits substantial residual shift such that $n_{1}\mathbb{E}[f_{i}^{(1)}]<(\sqrt{1+n_{1}/n_{0}}-1)n_{0}\mathbb{E}[f_{i}^{(0)}]$, then incorporating this source study will lead to negative transfer.
@@
  • (Covariate Shift) Our framework can relax the elusive conditions about the covariance matrices across studies (Assumption 4 in [tian2022transfer], Condition 4 in [zhang2022transfer], Assumption 4 in [huang2023estimation]). Specifically, our results neither assume the similarity of $\boldsymbol{\Sigma}^{(k)}$, nor involve the factor $C_{\boldsymbol{\Sigma}}$ mentioned above, hence achieving better performance under covariate shift.
@@
Under Conditions 1-3 and assumptions in Theorem 1, denote by $\widehat{\boldsymbol{\beta}}_{ora}$ the estimator obtained from (3) with the oracle transferable set $\mathcal{C}_{h}$. Then for any constant $\varepsilon>0$,
$\|\widehat{\boldsymbol{\beta}}_{ora}-\boldsymbol{\beta^{*}}\|_{2}\lesssim\sqrt{\frac{s\log p}{n_{\mathcal{C}_{h}}}}\frac{n_{\mathcal{C}_{h}}h_{2}}{|\mathcal{C}_{h}|n_{0}}+\sqrt{h_{1}}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\frac{n_{\mathcal{C}_{h}}h_{2}}{|\mathcal{C}_{h}|n_{0}}}.$
with probability at least $1-\varepsilon-(|\mathcal{C}_{h}|+1)A_{1}\exp(-A_{2}\log p)$.
As shown in Corollary 1, the terms $h_{1}$ and $h_{2}$ regulate the degree of parameter and residual shift, respectively, striking a balance between performance improvement and the risk of negative transfer. For the first term, a sufficient condition to avoid negative transfer, i.e, sharper than $\sqrt{s\log p/n_{0}}$, is to ensure $h_{2}=o(|\mathcal{C}_{h}|\sqrt{n_{0}/n_{\mathcal{C}_{h}}})$, which indicates that the information quality $n_{k}\mathbb{E}[f_{i}^{(k)}]$ in the oracle transferable set should exceed the order $\sqrt{n_{\mathcal{C}_{h}}n_{0}}\mathbb{E}[f_{i}^{(0)}]/|\mathcal{C}_{h}|$. The second term highlights the nuanced relationship between the shift levels $h_{1}$ and $h_{2}$, with a sufficient condition to avoid negative transfer being $h_{1}=o(s\sqrt{\log p/n_{0}}),h_{2}=o(|\mathcal{C}_{h}|\sqrt{n_{0}/n_{\mathcal{C}_{h}}})$. This generalizes the previous result [li2023estimation], allowing for better error control by managing both parameter and residual shift. To further clarify why incorporating $\tau$-th densities into the upper error bound is meaningful, we provide the following lower bound of $\ell_{2}$-estimation error in the minimax sense.
Condition 4.
Assume the second derivative of the logarithm of the conditional density functions $f^{(k)}(\cdot|\boldsymbol{x})$ exist and are uniformly bounded, i.e, $|[\log f_{i}^{(k)}(\cdot|\boldsymbol{x})]^{\prime\prime}|\leq L$ for any $\boldsymbol{x}$ and $k\in\{0\}\cup\mathcal{C}_{h_{1}}$ with the positive constant $L$.
Remark 4.
Condition 4 imposes a mild smoothness requirement on the conditional density function, which can be satisfied by many common distributions such as Gaussian, Logistic, and $t$ distributions.
Theorem 2.
Assume Condition 2 and 4 hold. Suppose the parameter space
$\boldsymbol{\Theta}(s,h_{1})=\{\boldsymbol{B}=(\boldsymbol{\beta},\boldsymbol{\delta}^{(1)},\dots,\boldsymbol{\delta}^{(K)}):\|\boldsymbol{\beta}\|_{0}\leq s,\max_{k\in\mathcal{C}_{h_{1}}}\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1}\}.$
If $\max\{s\log p/n_{\mathcal{C}_{h_{1}}},h_{1}(\log p/n_{0})^{1/2}\}=o(1)$, then
$\inf_{\widehat{\boldsymbol{\beta}}}\sup_{\boldsymbol{B}\in\boldsymbol{\Theta}(s,h_{1})}\mathbb{P}\left(\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}\|_{2}\gtrsim\sqrt{\frac{s\log p}{n_{\mathcal{C}_{h_{1}}}}}+\sqrt{\frac{s\log p}{n_{0}}}\wedge\left(\frac{h_{1}^{2}\log p}{n_{0}}\right)^{1/4}\wedge h_{1}\right)\geq\frac{1}{2},$
where $\widehat{\boldsymbol{\beta}}$ is an arbitrary estimator obtained from the data $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k\in\{0\}\cup\mathcal{C}_{h_{1}}}$.
To understand the lower bound in Theorem 2, we highlight two points. First, when the residual shift is ignored and the transfer procedure (3) is applied to the transferable set $\mathcal{C}_{h_{1}}$, the $l_{2}$-error bound $\sqrt{s\log p/n_{\mathcal{C}_{h_{1}}}}+\sqrt{h_{1}}(\log p/n_{0})^{1/4}$ established in Corollary 1 is nearly minimax-optimal. The first term attains the minimax-optimal rate, whereas the second term becomes slower when $h_{1}=o(\log p/n_{0})$ or $h_{1}\gg s\sqrt{\log p/n_{0}}$, reflecting the limitations of current transfer theory for high-dimensional quantile regression. Second, since the upper bounds are essentially sharp, when the parameter shift within the transferable set is already well controlled (e.g., $h_{1}=O(\sqrt{\log p/n_{0}})$), it remains worthwhile to improve estimation efficiency by further controlling the residual shift. This improvement is particularly meaningful when the sample sizes of source studies are moderate, which underscores the necessity of introducing the additional transferability measure $h_{2}$ in (2).